OpenStudy (anonymous):

Y=-4/7x-5, find the y coordinates of points (-7),(0),and (7) Please show work

OpenStudy (callisto):

If the given numbers are the x coordinates. Sub. the values given one by one into the given equation to get the respective value of y. Take -7 as an example Y=-4/7x-5 =-4/7 (-7)-5 = 4 -5 = -1 Can you do the rest?

OpenStudy (anonymous):

im still lost

OpenStudy (callisto):

in where?

OpenStudy (anonymous):

i got a fraction when i did it for the zero

OpenStudy (callisto):

Can you show your workings?

OpenStudy (anonymous):

y=-4/7x-5 =-4/7(0)-5 =-4/7-5 =4 3/7

OpenStudy (anonymous):

i got -9 when i did it for the 7

OpenStudy (callisto):

-4/7 (0) = ...?

OpenStudy (callisto):

For 7, you did it right :)

OpenStudy (anonymous):

-4/7(0)=0

OpenStudy (callisto):

Yes. So, y=-4/7x-5 =-4/7(0)-5 =0-5 =...

OpenStudy (anonymous):

-5

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

thank you :)

OpenStudy (callisto):

welcome!